# L-WISE: Boosting human image category learning through model-based image selection and enhancement

### v1

@@ 6 Methods
Predicting Image Difficulty. To predict the relative difficulty $d\in[0,1]$ of each image, we extract the logit value corresponding to the image’s ground truth class ($L_{\text{gt}}$) immediately upstream of the final softmax function. We sort the logits in descending order such that $L_{(1)}\geq L_{(2)}\geq...\geq L_{(n_{c_{i},s})}$. $n_{c_{i},s}$ is the number of images for a given class $c_{i}$ and training/validation/testing category designation $s$. We calculate difficulty percentile $d_{j}$ for image $j$ using the equation $d_{j}=\text{rank}(L_{(j)})/n_{c_{i},s}$.
@@
$x^{\prime}=x+\operatorname*{arg\,max}_{\|\delta\|<\epsilon}L_{\text{gt}}(x+\delta)-\frac{\alpha}{c-1}\sum_{c_{i}\in C,c_{i}\neq\text{gt}}L_{c_{i}}(x+\delta)$ (1)
@@
In equation 1, $\delta$ is a perturbation tensor of the same dimensionality as $x$ and an $\ell_{2}$ norm less than pixel budget $\epsilon$. $L_{\text{gt}}$ is the ANN’s logit score associated with the ground truth class, and $L_{c_{i}}$ is the logit associated with class $c_{i}$ ($c$ is the total number of classes, $C$ is the set of competing classes). $\alpha$ determines the extent to which logits for competing classes are minimized.We set $\alpha=0$ for the ImageNet animal classification experiments and $\alpha=1$ for the learning experiments.
@@
Image classification and learning experiments with human participants. We recruited 521 human subjects using the online platform Prolific. We allowed subjects to participate in multiple experiments, but only once for each of the three learning tasks. We used the JsPsych library (De Leeuw, 2015) with the JsPsychPsychophsics plugin (Kuroki, 2021) for all experiments. All experiments included 10-12.5% attention check trials where the participant is asked to classify an image of a circle or triangle. We analyzed data from participants with $\geq 90\%$ attention check accuracy.
@@
To measure the effects of enhancement on a presumably already-learned task, we tested the accuracy of human subjects at classifying 16 basic types of animals (frog, bird, dog, etc., see Appendix section S3). Images were shown for 17 milliseconds each, after which the participant was given 15 seconds to respond. All images were drawn from the validation set of ImageNet (Deng et al., 2009). Appendix Fig. S2 shows screenshots of the task as it appeared to participants. Each subject viewed interspersed images from 9 conditions: original images, images enhanced by maximizing $L_{\text{gt}}$ with $\epsilon=5,10,15,$ and $20$, images disrupted by minimizing $L_{\text{gt}}$ with $\epsilon=10$, and images enhanced with three baseline enhancement algorithms (see Fig. 1). Participants were notified after each incorrect response and given a small monetary bonus for each correct response (except for disrupted images). 62 participants viewed 18 images for each of 16 classes, 2 from each condition, in a shuffled ordering for a total of 288 trials per participant and 17,856 overall. These main trials followed a screening phase with 32 trials (200ms presentation times, 24 or more correct with at least one correct per class required to proceed, multiple screening attempts allowed) and a 32-trial warm-up phase with 17ms image presentations. Screening and warm-up phases used original, unmodified images, and data from these phases was not included in any analyses.
@@
The image category learning tasks consisted of either 4 (moths, dermoscopy) or 2 (histology) image classes. Participants were shown each image for up to 10 seconds and used the mouse (4-way tasks) or keyboard (binary task, “F” and “J” keys) to respond. After each trial, the participant was notified of the ground truth label and whether their response was correct. (screenshots in appendix Fig. S3). Each session consisted of 8 training blocks of 16 trials each (4 per class, or 2 per class for histology), and 2 testing blocks of 20 trials each during which no post-trial feedback was provided. Each block contained an equal number of images from each class in random order. Participants were informed upon recruitment that they could receive a progressively higher monetary bonus if their test-phase accuracy exceeded certain thresholds. Before participating in the main learning tasks, subjects had to first learn an easier binary classification task (leatherback vs. loggerhead turtles) and respond correctly to at least 7 of 8 test-phase trials. We randomly assigned ~30 participants per experimental condition, with ~60 participants for ablation study control conditions (min. 27, max. 68).
@@
Assisting learners with the L-WISE algorithm. L-WISE consists of two strategies applied in parallel: Enhancement Tapering (ET) and Difficulty Selection (DS). Both strategies operate only on images in the first 6 of 8 trial blocks in the training phase. In ET, we enhanced images in the first block of training-phase trials with $\epsilon=8$. $\epsilon$ is halved for each subsequent block until it is set to 0 after the 6th block. In DS, only images with $d<d_{b}$ were sampled for each block $b$. $d_{b}$ was incremented by 0.15 at the end of each of the first six blocks, beginning at $d_{1}=0.1$ and reaching $d_{7},d_{8}=1.0$. Determining an effective schedule of $\epsilon$ and $d_{\text{max}}$ did not require extensive hyperparameter tuning. After a pilot experiment in which we decreased $\epsilon$ linearly starting from $\epsilon=20$, (see appendix Fig. S9H-M), we switched to the $\epsilon$ schedule above changed no other hyperparameters for any of the three learning tasks/image domains. In the “shuffled” versions of DS and ET (Section 4.3), all trials in the first 6 blocks are selected or enhanced before a constrained shuffling procedure, where each image may switch positions with any other of the same class regardless of $d$ or $\epsilon$.
@@ 7 Ethics Statement
This study involved experiments with human participants conducted over the internet, using the Prolific platform for the main experiments and Amazon Mechanical Turk for pilot experiments. We followed a study protocol approved by the ((anonymized IRB information)). Participants provided informed consent before participating in any experiments. The experiments posed no greater than minimal risk to the participants. All participants were anonymous, and all data is de-identified. Participants were provided with our contact information, and that of the ((anonymized research office of our institution)), for any questions or concerns about the study. We calibrated the participant compensation amounts for each experiment to meet or exceed the equivalent of \$15.00 USD per hour, including during screening tasks. Participants were recruited using the "Standard Sample" option in Prolific, and were diverse in gender, age, and race/ethnicity (please see Table S2 for a demographic breakdown).
We hope that our work will eventually lead to practical and beneficial applications in education - for example, in the training of doctors in certain specialties such as pathology, radiology and dermatology where visual perceptual learning is an important part of clinical training. We wish to emphasize, however, that more work is needed before our methods can be safely applied in sensitive or high-stakes settings. For example, we apply our approach to improve human performance on a dermoscopy skin lesion classification task derived from the HAM10000 dataset (Tschandl et al., 2018). This dataset is heavily skewed towards images of pale skin, likely a reflection of the lower incidence of skin cancers such as melanoma among people with darker skin tones (Cormier et al., 2006). Models trained on this dataset are known to perform poorly for patients with darker skin (Daneshjou et al., 2022), where melanoma tends to have a different appearance, unfamiliarity with which on the part of clinicians contributes to delayed diagnosis and increased mortality among such patients (Thompson et al., 2023). It is plausible that maximizing the melanoma-associated logit of a robustified model perturbs the images to look more like an average presentation of melanoma (i.e., on light skin), which would risk imparting this bias onto the learner. A similar risk might apply to image selection: images with the highest groundtruth logits (which L-WISE presents at the beginning of learning) might tend to belong to specific subclasses or limited regions of the overall class distribution. The possibility of biased perturbations or image selections must be thoroughly investigated in future work before any applications of our methods in this domain.
@@ 8 Reproducibility Statement
We list all hyperparameters for training robustified neural networks and using them to enhance images in the appendix. Upon publication, we will provide source code that reproduces all data processing steps, figure plotting, and experiments. This includes experiments with human participants: we have developed a flexible framework for automated deployment of web-based image category learning experiments to a suite of cloud-based services, including hosting the task as an interactive web page and mechanisms for random group assignment and data collection. The experiments with human participants can therefore be reproduced and readily extended or modified with only a modest amount of configuration required.
@@ Acknowledgments
This work was supported in part by Harvard Medical School under the Dean’s Innovation Award for the Use of Artificial Intelligence, in part by Massachusetts Institute of Technology through the David and Beatrice Yamron Fellowship, in part by the National Institute of General Medical Sciences under Award T32GM144273, in part by the National Institutes of Health under Grant R01EY026025, and in part by the National Science Foundation under Grant CCF-1231216. The content is solely the responsibility of the authors and does not necessarily represent the official views of any of the above organizations. The authors would like to thank Andrei Barbu, Roy Ganz, Katherine Harvey, Michael J. Lee, Richard N. Mitchell, and Luke Rosedahl for sharing their helpful insights into our work at various times.
@@
The ResNet-50 model adversarially trained on ImageNet was used directly to generate perturbations and difficulty rankings for the 16-way animal classification task, using the logits of the original, fine-grained ImageNet classes (i.e., not the 16 superclasses, “grasshopper” not “insect”) for both enhancement and difficulty prediction. The same model was adversarially fine-tuned on the HAM10000 and MHIST datasets before their application (as part of L-WISE) to the dermoscopy and histology tasks respectively. Generally, we trained the models on all available classes in each dataset. For example, we fine-tuned on all 7 classes of the HAM10000 dermoscopy dataset (Tschandl et al., 2018), even though we only used 4 of them in the learning tasks. When enhancing the images, we include only classes that are part of the experimental tasks as competing classes to have their logits minimized (see main-text equation 1).
@@
For the moth task, we adversarially fine-tuned the (adversarially) iNaturalist-pretrained model on the four moth classes to be used in the task (which are part of iNaturalist). We subjectively judged the perturbations from this fine-tuned model to be more compelling than those generated using the iNaturalist-pretrained model without fine-tuning, as the four classes of interest are among the 10,000 iNaturalist classes.
All fine-tuning used $\epsilon=1.0$ for adversarial training. We tuned various hyperparameters (learning rate, data augmentation strategies, etc.) during fine-tuning to maximize validation set performance on the images to be used for the learning tasks. We fine-tuned the entire network end-to-end for each task.
@@
Our choice of $\epsilon$=3 for ImageNet pretraining follows Gaziv et al. (2024), who found this to be an optimal choice for perturbations that disrupt category perception (relative to $\epsilon$=1 and $\epsilon=10$). In practice, we found that the models were unable to learn finer-grained tasks with training-time adversarial perturbations as large as $\epsilon$=3 (iNaturalist pretraining and fine-tuning on moth photos, dermoscopy images, and histology images) - therefore, we reverted to $\epsilon$=1 for these settings.
@@
To enhance images in a category-specific manner, we perform the optimization of Equation 1 (main text) in a series of steps using projected gradient descent (equation 2), where $k$ denotes the optimization step, $\eta$ the step size, and $\text{Proj}_{\epsilon}$ a projection onto a hypersphere of radius $\epsilon$ with $x$ at its center. All enhancements with pixel budget $\epsilon$ use $\text{ceil}(2\epsilon)$ steps of $\eta=0.5$ in a $224\times 224\times 3$ pixel space.
@@
$\delta_{k+1}=\text{Proj}_{\epsilon}\bigg{(}\delta_{k}-\eta\nabla_{\delta}\big{(}L_{\text{gt}}(x+\delta_{k})-\frac{\alpha}{c-1}\sum_{c_{i}\in C,c_{i}\neq\text{gt}}L_{c_{i}}(x+\delta_{k})\big{)}\bigg{)}$ (2)
@@
Fig. S1 shows several example images enhanced with $\epsilon=20$ using this approach by $\epsilon=3$ adversarially-pretrained ResNet-50, along with difference images and heat maps produced by the same method as Figs. 3-4 in the main text.
@@
All images presented in all experiments were of size $224\times 224\times 3$, matching the input dimensions of ResNet-50. Before presentation or any model-based enhancement or difficulty prediction, original images were resized such that the shortest dimension (width or height) was 224 pixels, and then center-cropped to $224\times 224$. Any single-channel grayscale images were converted to RGB before further processing. The baseline enhancement algorithms Contrast-Limited Adaptive Histogram Equalization (CLAHE (Zuiderveld, 1994)), Multi-Scale Retinex with Color Restoration (MSRCR (Jobson et al., 1997; Petro et al., 2014)), and the “Auto” image tuning feature in Adobe Photoshop Lightroom (Auto-LR (Adobe Inc., 2024)) were applied before the resizing and center-cropping operations.
@@
Figure S2: Task interface for ImageNet animal classification with human participants. Subjects classified images among 16 categories. During each trial, the subject clicks the fixation cross (panel A) and the image is flashed for 17 milliseconds (panel B) with a 200ms presentation of a blank screen immediately before and after. The mouse cursor is hidden during the image presentation. Images are presented such that they subtend approximately 6 degrees of visual angle, with calibration for each participant using a blind-spot calibration procedure. After viewing the image, the participant clicks one of 16 buttons shown in panel C, which are randomly rotated in position every trial, within 15 seconds. For incorrect responses (except on images that are disrupted via ground truth logit minimization), or if 15 seconds elapses without a response, the participant is shown the black X for one second (panel D). Otherwise, no explicit feedback is given and the next trial begins immediately. Attention check trials featuring an image of a circle or triangle (see Fig. S3D for an example image) were interspersed with the main trials. For the attention check trials, two of the animal icons in panel C were randomly selected to be replaced with circle and triangle icons.
@@
We curated 16 sets of ImageNet classes corresponding to 16 basic animal superclasses for our basic animal classification experiments (Figs. 1-S2), adapting and expanding the Restricted ImageNet dataset defined in the Robustness library (Engstrom et al., 2019). The assignment of specific animal classes to each superclass is listed below:
@@
We excluded certain classes on a case-by-case basis in an attempt to minimize errors due to misunderstanding the animal categories as opposed to errors of visual perception. For example, we did not include porcupines, hedgehogs, or beavers in the “rodent” class (as many people may not recognize these as rodents), and we do not include eels in the “fish” class due to the possibility of confusion with snakes. Our classification of rabbits and hares as “rodents” is technically incorrect, as they were reclassified to the order Lagomorpha in 1912 Chapman & Flux (2008) (we thank the participant who notified us of this).
We mistakenly classified rabbits and hares as “rodents,” although they were reclassified to the order Lagomorpha in 1912.
@@
Figs. S2 and S3 show the task interfaces for the 16-way animal classification and 4-way image category learning experiments with human participants, respectively. The positions of the four buttons used to indicate responses are randomly permuted for each participant.
To minimize biases stemming from prior declarative knowledge of the tasks or classes, for each participant in the 4-category learning tasks we randomly assign a four-letter, two-syllable name from the set "Ajax," "Eris," "Leda," and "Tyro" drawn from Greek mythology, each having four letters, two syllables, two consonants, and two phonetic vowels. We found no evidence that associating certain categories with certain aliases consistently affected test-phase accuracy (see Figs. S12 and S13).
We used a different approach for the binary histology task that employed the MHIST dataset (Wei et al., 2021), giving benign hyperplostic polyp the alias “benign” and sessile serrated adenoma the alias “malignant” (although sessile serrated adenoma is really a pre-cancerous legion). The histology task has an interface very similar in appearance to that of the 4-way tasks (as shown in Fig. S3), except that the “benign” and “malignant” buttons always appear either on side of the presented image (in a random order for each participant), and the participant responds by pressing the F key for the left-hand category or J for the right instead of clicking one of the buttons.

### v2

@@ 6 Methods
Predicting Image Difficulty. To predict the relative difficulty $d\in[0,1]$ of each image, we extract the logit value corresponding to the image’s ground truth class ($L_{\text{gt}}$) immediately upstream of the final softmax function. We sort the logits in descending order such that $L_{(1)}\geq L_{(2)}\geq...\geq L_{(n_{c,s})}$. $n_{c,s}$ is the number of images for a given class $c$ and training/validation/testing split designation $s$. We calculate class-specific difficulty percentile $d_{j}$ for image $j$ using the equation $d_{j}=\text{rank}(L_{(j)})/n_{c,s}$.
@@
$x^{\prime}=x+\operatorname*{arg\,max}_{\|\delta\|<\epsilon}\big{(}L_{\text{gt}}(x+\delta)-\frac{\alpha}{\lvert C\rvert-1}\sum_{c\in C:c\neq\text{gt}}L_{c}(x+\delta)\big{)}$ (1)
@@
In Equation 1, $\delta$ is a perturbation tensor of the same dimensionality as $x$ and with an $\ell_{2}$ norm less than pixel budget $\epsilon$. $L_{\text{gt}}$ is the ANN’s logit score associated with the ground truth class, and $L_{c}$ is the logit associated with class $c$ ($C$ is the set of all classes, with cardinality $\lvert C\rvert$). $\alpha$ determines the extent to which logits for competing classes are minimized.We set $\alpha=0$ for the ImageNet animal classification experiments and $\alpha=1$ for the three fine-grained image category learning experiments.
@@
Image classification and learning experiments with human participants. We recruited 521 human subjects using the online platform Prolific. We allowed subjects to participate in multiple experiments, but only once for each of the three learning tasks. We used the JsPsych library (De Leeuw, 2015) with the JsPsychPsychophsics plugin (Kuroki, 2021) for all experiments. All experiments included 10-12.5% attention check trials where the subject classifies an image of a circle or triangle (e.g., see Appendix Fig. S4). We analyzed data from subjects with $\geq 90\%$ attention check accuracy.
@@
To measure the effects of enhancement on a presumably already-learned task, we tested the accuracy of human subjects at classifying 16 basic types of animals (frog, bird, dog, etc., see Appendix section S3). Images were shown for 17 milliseconds each, after which the participant was given 15 seconds to respond. All images were drawn from the validation set of ImageNet (Deng et al., 2009). Appendix Fig. S3 shows screenshots of the task as it appeared to participants. In our main experiment with this task (Fig. 1), each subject viewed interspersed images from 9 conditions: original images, images enhanced by maximizing $L_{\text{gt}}$ with $\epsilon=5,10,15,$ and $20$, images enhanced with one of three off-the-shelf baseline methods, and images disrupted by minimizing $L_{\text{gt}}$ (internal control). Participants were notified after each incorrect response and given a small monetary bonus for each correct response (except for disrupted images). 62 participants viewed 18 images for each of 16 classes, 2 from each condition, in a shuffled ordering for a total of 288 trials per participant and 17,856 overall. These main trials followed a screening phase with 32 trials (200ms presentation times, 24 or more correct with at least one correct per class required to proceed, multiple screening attempts allowed) and a 32-trial warm-up phase with 17ms image presentations. Screening and warm-up phases used original, unmodified images, and data from these phases were not included in any analyses.
@@
The image category learning tasks consisted of either 4 (moths, dermoscopy) or 2 (histology) image classes. Participants were shown each image for up to 10 seconds and used the mouse (4-way tasks) or keyboard (binary task, “F” and “J” keys) to respond. After each trial, the participant was notified of the ground truth label and whether their response was correct (screenshots in Appendix Fig. S4). Each session consisted of 8 training blocks of 16 trials each (4 per class, or 2 per class for histology), and 2 testing blocks of 20 trials each during which no post-trial feedback was provided. Each block contained an equal number of images from each class in a random order. Participants were informed upon recruitment that they could receive a progressively higher monetary bonus if their test-phase accuracy exceeded certain thresholds. Before participating in the main learning tasks, subjects had to first learn an easier binary classification task (leatherback vs. loggerhead turtles) and respond correctly to at least 7 of 8 test-phase trials. We randomly assigned ~30 participants per experimental condition, except the ablation study control groups of ~60 participants (overall min. 27, max. 68).
@@
Assisting learners with the L-WISE algorithm. L-WISE consists of two strategies applied in parallel: Enhancement Tapering (ET) and Difficulty Selection (DS). Both strategies operate only on images in the first 6 of 8 trial blocks in the training phase. In ET, we enhanced images in the first block of training-phase trials with $\epsilon=8$. $\epsilon$ is halved for each subsequent block until it is set to 0 after the 6th block. In DS, only images with $d<d_{b}$ were sampled for each block $b$. $d_{b}$ was incremented by 0.15 at the end of each of the first six blocks, beginning at $d_{1}=0.1$ and reaching $d_{7},d_{8}=1.0$. Determining an effective schedule of $\epsilon$ and $d_{b}$ did not require extensive hyperparameter tuning. After a pilot experiment in which we decreased $\epsilon$ linearly starting from $\epsilon=20$, (see Appendix Fig. S12H-M), we switched to the $\epsilon$ schedule above and changed no other hyperparameters for any of the three learning tasks/image domains. In the “shuffled” versions of DS and ET (Section 4.3), images in blocks 1-6 are selected or enhanced before a constrained shuffling procedure, whereby each image in blocks 1-6 may switch positions with any other of the same class regardless of $d$ or $\epsilon$.
@@ 7 Ethics Statement
This study involved experiments with human participants conducted over the internet, using the Prolific platform for the main experiments and Amazon Mechanical Turk for pilot experiments. We followed a study protocol approved by Boston Children’s Hospital’s Institutional Review Board. Participants provided informed consent before participating in any experiments. The experiments posed no greater than minimal risk to the participants. All participants were anonymous, and all data is de-identified. Participants were provided with our contact information, and that of the Office of Clinical Investigations at Boston Children’s Hospital, for any questions or concerns about the study. We calibrated the participant compensation amounts for each experiment to meet or exceed the equivalent of \$15.00 USD per hour, including during screening tasks. Participants were recruited using the “Standard Sample” option in Prolific, and were diverse in gender, age, and race/ethnicity (please see Table S2 for a demographic breakdown).
We hope that our work will eventually lead to practical and beneficial applications in education - for example, in the training of doctors in specialties such as pathology, radiology and dermatology where visual perceptual learning is particularly important. We wish to emphasize, however, that more work is needed before our methods can be safely applied in sensitive or high-stakes settings. For example, we apply our approach to improve human performance on a dermoscopy skin lesion classification task derived from the HAM10000 dataset (Tschandl et al., 2018). This dataset is heavily skewed towards images of pale skin, likely a reflection of the lower incidence of skin cancers such as melanoma among people with darker skin tones (Cormier et al., 2006). Models trained on this dataset are known to perform poorly for patients with darker skin (Daneshjou et al., 2022), where melanoma tends to have a different appearance, unfamiliarity with which on the part of clinicians contributes to delayed diagnosis and increased mortality among such patients (Thompson et al., 2023). It is plausible that maximizing the melanoma-associated logit of a robustified model perturbs the images to look more like an average presentation of melanoma (i.e., on light skin), which would risk imparting this bias onto the learner. A similar risk might apply to image selection: images with the highest ground truth logits (which L-WISE presents at the beginning of learning) might tend to belong to specific subclasses or limited regions of the overall class distribution. The possibility of biased perturbations or image selections must be thoroughly investigated in future work before any applications of our methods in this domain.
@@ 8 Reproducibility Statement
We list all hyperparameters for training robustified neural networks and using them to enhance images in the Appendix. We provide source code that enables reproduction of all data processing steps, experiments, statistical analyses, and figure plotting. This includes experiments with human participants: we have developed a flexible framework for automated deployment of web-based image category learning experiments, including hosting the task as an interactive web page and cloud-based mechanisms for random group assignment and data collection. The experiments with human participants can therefore be readily reproduced and extended or modified with only a modest amount of configuration required.
@@ Acknowledgments
This work was supported in part by Harvard Medical School under the Dean’s Innovation Award for the Use of Artificial Intelligence, in part by Massachusetts Institute of Technology through the David and Beatrice Yamron Fellowship, in part by the National Institute of General Medical Sciences under Award T32GM144273, in part by the National Institutes of Health under Grant R01EY026025, and in part by the National Science Foundation under Grant CCF-1231216. The content is solely the responsibility of the authors and does not necessarily represent the official views of any of the above organizations. The authors would like to thank Yousif Kashef Al-Ghetaa, Andrei Barbu, Pavlo Bulanchuk, Roy Ganz, Katherine Harvey, Michael J. Lee, Richard N. Mitchell, and Luke Rosedahl for sharing their helpful insights into our work at various times.
@@
The ResNet-50 model adversarially trained on ImageNet was used directly to generate perturbations and difficulty rankings for the 16-way animal classification task, using the logits of the original, fine-grained ImageNet classes (i.e., not the 16 superclasses, “grasshopper” not “insect”) for both enhancement and difficulty prediction. The same model was adversarially fine-tuned on the HAM10000 and MHIST datasets before their application (as part of L-WISE) to the dermoscopy and histology tasks respectively. Generally, we trained the models on all available classes in each dataset. For example, we fine-tuned on all 7 classes of the HAM10000 dermoscopy dataset (Tschandl et al., 2018), even though we only used 4 of them in the learning task for humans. When enhancing the images, we include only classes that are part of the experimental tasks as competing classes to have their logits minimized (see main-text Equation 1).
@@
For the moth task, we adversarially fine-tuned the (adversarially) iNaturalist-pretrained model on the four moth classes to be used in the task. We subjectively judged the perturbations from this fine-tuned model to be of higher quality than those generated using the iNaturalist-pretrained model without fine-tuning, as the four classes of interest are among the 10,000 iNaturalist classes. All adversarial fine-tuning used $\epsilon=1.0$ with 7 gradient steps of size 0.3. We used learning rates of 0.0001, 0.0004, and 0.001, and batch sizes of 32, 64, and 16, for Idaea moth, dermoscopy, and histology fine-tuning respectively. We fine-tuned the entire network end-to-end for each task.
@@
Our choice of $\epsilon=3$ for ImageNet pretraining follows Gaziv et al. (2023), who found this to be an optimal choice for generating perturbations that disrupt category perception (relative to $\epsilon=1$ and $\epsilon=10$). In practice, we found that the models were unable to learn finer-grained tasks with training-time adversarial perturbations as large as $\epsilon=3$ (iNaturalist pretraining, and fine-tuning on moth photos, dermoscopy images, and histology images) - therefore, we reverted to $\epsilon=1$ for these settings.
@@
To enhance images in a category-specific manner, we perform the optimization of Equation 1 (main text) in a series of steps using projected gradient ascent (Equation 2), where $k$ denotes the optimization step, $\eta$ the step size, and $\text{Proj}_{\epsilon}$ a projection onto a hypersphere of radius $\epsilon$ with original image $x$ at its center (see text below Equation 1 for definitions of other symbols).
@@
$\delta_{k+1}=\text{Proj}_{\epsilon}\bigg{(}\delta_{k}-\eta\nabla_{\delta}\big{(}L_{\text{gt}}(x+\delta_{k})-\frac{\alpha}{\lvert C\rvert-1}\sum_{c\in C:c\neq\text{gt}}L_{c}(x+\delta_{k})\big{)}\bigg{)}$ (2)
@@
Fig. S1 shows several example images enhanced with $\epsilon=20$ using logit maximization by adversarially-pretrained ResNet-50 ($\epsilon=3$), along with difference images and heat maps produced by the same method as Figs. 3-4 in the main text. Throughout our experiments, robustified ResNet-50 enhancements with pixel budget $\epsilon$ used $\text{ceil}(2\epsilon)$ steps of $\eta=0.5$ in a $224\times 224\times 3$ pixel space. This formula seems somewhat model-dependent and sometimes requires adjustment: for example, when enhancing images with robustified XCiT (see Figs. S9-S10), $\text{ceil}(4\epsilon)$ steps were required instead of $\text{ceil}(2\epsilon)$ to reach a similar effective perturbation size.
We generate the heat maps in Figs. 3, 4, and S1 by subtracting the enhanced image $x^{\prime}$ from the original image $x$ element-wise: $\delta=x^{\prime}-x$. We calculate the magnitude of the changes as $\delta^{2}$. We apply smoothing using 2D convolution, and normalize the result to have all values between 0 and 1, to produce $\delta_{\text{norm}}$. We then produce the heat map by setting the red channel to $255\times\delta_{\text{norm}}$ and the blue channel to $255\times(1-\delta_{\text{norm}})$. The resulting image shows red in regions where larger changes have taken place, and blue in regions where smaller or no changes have taken place. In Figs. 34, and S1, we superimpose a translucent version of the heat maps ($\alpha=0.7$) over the original images. Averaging the heat maps across all ImageNet validation set images (Fig. S2) indicates that changes to the images tend to occur in the central regions of the images more than in the periphery, consistent with the observation that our image enhancement approach tends to primarily change image regions corresponding to the main subject of each image.
Figure S2: Pixel-value changes during enhancement of ImageNet images are biased towards the center of the image. This heat map indicates which spatial regions of the ImageNet animal images were changed the most on average during logit-maximization enhancement with $\epsilon=20$. Regions that were changed more on average are more red, and regions that were changed less on average are more blue. The heat map was generated by averaging the normalized absolute pixel value changes across all 2400 ImageNet validation set images that we used for our 16-way animal classification experiments.
@@
All images presented in all psychophysics experiments were of size $224\times 224\times 3$, matching the input dimensions of ResNet-50. Before presentation or any model-based enhancement or difficulty prediction, original images were resized such that the shortest dimension (width or height) was 224 pixels, and then center-cropped to $224\times 224$. Any single-channel grayscale images were converted to RGB before further processing. The baseline enhancement algorithms Contrast-Limited Adaptive Histogram Equalization (CLAHE (Zuiderveld, 1994)), Multi-Scale Retinex with Color Restoration (MSRCR (Jobson et al., 1997; Petro et al., 2014)), and the “Auto” image tuning feature in Adobe Photoshop Lightroom (Auto-LR (Adobe Inc., 2024)) were applied before the resizing and center-cropping operations.
@@
Figure S3: Task interface for ImageNet animal classification with human participants. Subjects classified images among 16 categories. During each trial, the subject clicks the fixation cross (panel A) and the image is displayed for 17 milliseconds (panel B) with a 200ms presentation of a blank screen immediately before and after. The mouse cursor is hidden during the image presentation. Images are presented such that they subtend approximately 6 degrees of visual angle, with calibration for each participant using a blind-spot calibration procedure. After viewing the image, the participant clicks one of the 16 buttons shown in panel C, which are randomly rotated in position every trial, within a 15-second time limit. For incorrect responses, or if 15 seconds elapses without a response, the participant is shown the black X for one second (panel D). Otherwise, no explicit feedback is given and the next trial begins immediately. Attention check trials featuring an image of a circle or triangle (see Fig. S4D for an example image) were interspersed with the main trials. For the attention check trials, two of the animal icons in panel C were randomly selected to be replaced with circle and triangle icons.
@@
We curated 16 sets of ImageNet classes corresponding to 16 basic animal superclasses for our basic animal classification experiments (e.g., see Figs. 1 and S3), adapting and expanding the Restricted ImageNet dataset defined in the Robustness library (Engstrom et al., 2019). The assignment of specific animal classes to each superclass is listed below:
@@
We excluded certain classes on a case-by-case basis in an attempt to minimize errors due to misunderstanding the animal categories, as opposed to errors of visual perception. For example, we did not include porcupines or beavers in the “rodent” class (as many people may not recognize these as rodents), and we did not include eels in the “fish” class due to the possibility of confusion with snakes. We mistakenly classified rabbits and hares as “rodents” given that they were reclassified to the order Lagomorpha in 1912 Chapman & Flux (2008) (we thank the participant who notified us of this).
@@
Figs. S3 and S4 show the task interfaces for the 16-way animal classification and 4-way image category learning experiments with human participants, respectively. For the learning experiments, the positions of the four buttons used to indicate responses are randomly permuted for each participant.
To minimize biases stemming from any priors induced by the class names, for each participant in the 4-category learning tasks we randomly assign a four-letter, two-syllable alias from the set “Ajax,” “Eris,” “Leda,” and “Tyro.” These names are drawn from Greek mythology, and each has four letters, two syllables, two consonants, and two phonetic vowels. We found no evidence that associating certain categories with certain aliases consistently affected test-phase accuracy (see Figs. S15 and S16).
We used a different approach for the binary histology task that employed the MHIST dataset (Wei et al., 2021), giving “benign hyperplastic polyp” the alias “benign” and sessile serrated adenoma the alias “malignant” (although sessile serrated adenoma is actually a pre-cancerous lesion). The histology task has an interface very similar in appearance to that of the 4-way tasks (as shown in Fig. S4), except that the “benign” and “malignant” buttons always appear on either side of the presented image (in a random order for each participant), and the participant responds by pressing the F key for the left-hand category or J for the right instead of clicking one of the buttons.
@@
Figure S5: The observed relationship between robust model ground truth logits and human error rates is not sensitive to trial inclusion criteria. For our main analysis regarding predicting ImageNet animal image recognition difficulty illustrated in Fig. 1A, we included all trials with original images but also trials featuring images modified by off-the-shelf control enhancement methods (Adobe Lightroom, CLAHE, and MSRCR), and trials with image perturbations from non-adversarially-trained ANNs (i.e., Vanilla ResNet-50 and CutMix ResNet-50 from Fig. S7). Panel A above reproduces Fig. 1A1, while panel B replicates the same analysis but strictly including only trials with natural, unmodified images. The numbers above each point indicate how many image trial observations were included in the corresponding bin (the size of the vertical 95% confidence intervals is sensitive to this). AUC=Area Under the Receiver Operating Characteristic Curve, p-values derived from Wald statistic on the logistic regression coefficient for the ground truth logit predictor.